# Parallel Batch Scheduling With Incompatible Job Families Via Constraint Programming

### v3

@@ I Introduction
Existing exact methods in the literature for p-batching include mixed-integer programming (MIP) [6, 18] and Constraint Programming (CP) models [8, 18]. MIP models do not scale well and are outperformed by the CP approaches [2, 18]. However, the existing CP model for p-batching with incompatible jobs is solver-dependent as it relies on specific functions and global constraints only available in a well know commercial CP solver. This paper expands the body of knowledge around this problem by proposing four new CP models that can be implemented in other solvers: one model that relies on automaton constraints, available in different open-source solvers; and other three models integrate assignment and scheduling decisions with different strategies and global constraints, available in many commercial and open-source solvers. Furthermore, the paper conducts thorough computational experiments that demonstrate the implementation flexibility of the proposed models, and compares them under different objectives.
@@ II Literature Review
For $\beta$, the focus is on articles that deal with multiple incompatible families (i.e., $P|p-batch,incompatible|\gamma$), although key articles with a single compatible family (i.e., $P|p-batch|\gamma$) are also considered, as the paper extends some of their concepts to the incompatible case. The review also includes some articles that consider non-identical job sizes ($s_{j}$) and release times ($r_{j}$). For $\gamma$, the reviewed objective functions primarily fall into three categories: (i) utilization objectives such as minimizing the makespan ($C_{\max}$); (ii) cycle time objectives such as minimizing the total completion time (TCT) and its weighted counterpart (TWCT); and (iii) on-time delivery objectives such as minimizing earliness, lateness, tardiness, number of tardy jobs, including their weighted variants [2]. However, since the diffusion area is usually a bottleneck in the wafer fab, fab leaders often focus on two key aspects: the cycle time of the jobs to increase the effective movement of products [19] and the utilization of the diffusion furnaces. Thus, this paper focuses on minimizing the TWCT, which specifically targets the cycle time performance [2], and the makespan ($C_{\max}$), which specifically targets the utilization performance [2].
@@ VI-A The Automaton model
The Automaton (AU) model derives its name from the automaton global constraint it relies on, which is available in different open-source CP solvers [41, 22, 42]. Automaton constraints model finite state machines, enabling the representation of feasible transitions between states. In the AU model, these finite states represent the elapsed time when processing a batch of each family. Consider the automaton depicted in Figure 2, where nodes are states and arcs are transition possibilities. The descriptions in each state help understand the figure and the numbers in parenthesis ultimately identify them. Each arc is identified by a triplet $(i,k,\tau)$, where $i$ is the initial state, $k$ is the end state, and $\tau$ is the required transition value to trigger such transition from $i$ to $k$. For example, if a machine is empty and receives a transition value of 0, then the machine remains empty. But if it receives a transition value of 1, it means that a new batch of family 1 starts on the machine. From this point, the only possible transition values that the automaton can receive are the necessary to progress the states until completing the processing time ${\delta}^{1}$ of the batch. Once the batch is completed, three possible transitions can be made: either the machine becomes empty (0), or a new batch of family 1 starts (1), or a new batch of family 2 starts (${\delta}^{1}+1$). From that point, a new sequence of transition values must be given such that the state transitions respect the automaton structure.
@@ VI-B The Assign-and-Schedule Model
Objective function (3) minimizes the TWCT. Constraints (3) assign each job to exactly one possible batch. Constraints (3) deactivate the presence of the batches if no jobs are assigned to them. Constraints (3) activate the presence of the batches if at least one job is assigned to them. Constraints (3) ensure that if a batch is used, its size respects the maximum batch capacity. Constraints (3) guarantee that batches are scheduled once all the jobs in them are available. Constraints (3) schedules each batch on exactly one machine, if present. Finally, constraints (3) use the global constraint noOverlap, which ensures that the present interval variables in the provided sequence do not overlap. Hence, constraints (3) ensure that only one batch is processed at a time on each machine.
@@ VI-C The Synchronized Model
Objective function (4) minimizes the TWCT. Constraints (4) assign each job to exactly one batch on exactly one machine. Constraints (4) allow each batch to be processed in at most one machine. Constraints (4) use the synchronize global constraint, which takes two inputs: an interval variable $v=x_{jm}^{b}$ and a set of interval variables $B=\set{y_{bm}}$. If $v$ is present, this constraint attempts to synchronize it with any intervals in $B$ that are also present [40], i.e., $\texttt{presenceOf}(v)=1\implies\texttt{startOf}(v)=\texttt{startOf}(b)\wedge\texttt{endOf}(v)=\texttt{endOf}(b),\forall\leavevmode\nobreak\ b\in B$ s.t. $\texttt{presenceOf}(b)=1$. In practice, when the interval variable representing the job being assigned to a batch on a machine (i.e., $x_{jm}^{b}$) is part of the solution, the synchronize constraint attempts to align its start and end times with those of the batch on the machine (i.e., $y_{bm}$), if it is also present. However, the synchronize constraint only ensures synchronization if both interval variables are present; it does not inherently enforce the presence of $y_{bm}$ when $x_{jm}^{b}$ is scheduled. This subtle issue, where the presence of $x_{jm}^{b}$ should logically imply the presence of $y_{bm}$ is addressed in the original model by Cakici et al. [10] by means of an additional constraints that defines the correct duration of the batches depending on the jobs assigned to it, if any. This constraint is necessary in their AM application as the duration of the batch depends on the size of the parts 3D-printed in it. However, this constraint is not required in the context of the diffusion area studied in this paper since the duration of the batches is already predefined. Nonetheless, the presence of the correct batches is still needed to be enforced. To resolve this, constraints (4) introduce an explicit implication constraint which ensures that, whenever $x_{jm}^{b}$ is included in the solution, the corresponding batch interval $y_{bm}$ is also present. Constraints (4) use the noOverlap global constraint to ensure that the present interval variables in the machine sequence do not overlap, guaranteeing at most one batch being processed at a time on each machine. Finally, constraints (4) guarantee that the load of the batches respect the maximum capacity.
@@ VII Computational experiments and results
and then setting $r_{j}\sim U\set{[C_{\max,lb}]}$. Finally, the maximum batch size for each family is generated as
$u_{f}\sim U(\{10,15,20,25,30,35,40,45,50,75,100,125,150\}).$
@@
Table II summarizes the models implemented in each solver. The A model was only implemented in CP Optimizer because its state functions, alwaysEqual global constraint, and alignment parameters are only available int. On the other hand, the AU model was only implemented in CP-SAT because the automaton global constraint required is not available in CP Optimizer (although it can be implemented in Choco-solver [22] and MiniZinc [42]). The AS, S, and RS models (and their variants with SB constraints) are implemented in both solvers.
@@
Figures 69 summarize our results. The graphs for TWCT focus on Figures 6 and 7, while Figures 8 and 9 detail the results for the $C_{\max}$ objective. Each graph presents the following characteristics:
@@
    Average gap (%) with its 95% confidence interval (CI).
@@
The gap for each instance is computed as follows. First, we obtain the minimum objective value $z_{i,obj}^{*}$ across all models and both solvers for each instance $i$ and objective $obj\in\{TWCT,C_{\max}\}$. Then, for a given model and solver, the gap is defined as $\frac{|z_{i,obj,\text{model, solver}}-z_{i,obj}^{*}|}{|z_{i,obj}^{*}|}$.
@@ VII-A Results with the TWCT objective
Figure 6(b) presents the average gap (%) along with its 95% confidence interval for each model in CP Optimizer. Although the A model did not prove optimality within the time limit in any instance, it consistently finds the best solutions, evident by the constant 0% gap of the black series. In contrast, the AS and AS+SB models exhibit the worst average gaps, with wider confidence intervals. This is partly due to their failure to find solutions in several instances, resulting in fewer data points for averaging. Notably, a series break for these models appears at 200 jobs, 6 families, and 10 machines, reflecting a complete failure to solve any instance in that class. Furthermore, the consistently lower gaps of the AS+SB model compared to the AS model underscore the beneficial impact of incorporating SB constraints in the search process.
@@
Figure 7 focus on the results of CP-SAT, which combines CP methods and satisfiability (SAT) methods in the search process. SAT methods benefit from binary structures in the problem. For this reason, in this solver, the AS model is able to find solutions on all the instances, as evidenced in Figure 7(a). The distinction between assignment and scheduling decisions of the AS model provides the proper structure for both CP and SAT methods to properly work, allowing the AS model to find solutions in all the instances; and even declare optimality in one of them. In contrast, the RS+SB model is unable to find solutions in most of the largest instances, precisely because of the additional SB constraints that become harder to satisfy within the time limit. The RS model is unable to find solutions in the largest instances with 200 job, precisely because of the amount of job-in-batch decisions that need to be made. On the other hand, the AU model clearly shows seasonal behavior around 10 machines.
When examining the average gaps in CP-SAT from Figure 7(b), it becomes evident that no model consistently matches the best results obtained by the A model in CP Optimizer. Nonetheless, the gaps are very close. The AU gap increases with the size of the instances. This is because of the $\tau_{mt}$ variables that represent the transition values provided to the automaton at every time step of the scheduling horizon, which increases with the number of jobs. Additionally, the element constraints set the appropriate transition values at the start times of the jobs on each machine. However, they do so only if the jobs are actually processed on such machine. While this could be seen as a reified constraint, it is not. Instead, this is indirectly achieved with the modified start times $\hat{s}_{jm}$ and modified transition values $\hat{\tau}_{jm}$. When the instances grow, more of these modified decisions have to be made, negatively impacting the performance of these element constraints, and causing the average gap increase.
@@
The best gaps in CP-SAT belong to the AS model, followed by the RS model, very close to 0%. Interestingly, the very same reason that makes The AS model model fall back in CP Optimizer, makes it shine in CP-SAT instead. The clear distinction between assignment and scheduling logic is well handled in CP-SAT thanks to the SAT component of this solver. Additionally, the RS model consistently produces comparable gaps of the AS model. However, as seen in Figure 7(a), the RS model is not able to find solutions in all the instances.
@@ VIII Concluding remarks
  • The A model consistently produces high-quality solutions (0% gap) but is limited by its reliance on proprietary solver features only available in the commercial solver CP Optimizer.
  • The clear separation between binary assignment and batch scheduling decisions enables the AS model to excel in the open-source solver CP-SAT. Thanks to the satisfiability components of the solver, which efficient handle binary structures, the AS model achieves very small gaps and even proves optimality in isolated cases.

### v4

@@ I Introduction
Existing exact methods in the literature for p-batching include mixed-integer programming (MIP) [6, 18] and Constraint Programming (CP) models [8, 18]. MIP models do not scale well and are outperformed by the CP approaches [2, 18]. However, the existing CP model for p-batching with incompatible jobs is solver-dependent as it relies on specific functions and global constraints only available in a well known commercial CP solver. This paper expands the body of knowledge around this problem by proposing four new CP models that can be implemented in other solvers: one model that relies on automaton constraints, available in different open-source solvers; and other three models integrate assignment and scheduling decisions with different strategies and global constraints, available in many commercial and open-source solvers. Furthermore, the paper conducts thorough computational experiments that demonstrate the implementation flexibility of the proposed models, and compares them under different objectives.
@@ II Literature Review
Bottlenecks in semiconductor manufacturing are dynamic and can occur in various areas depending on the process technology, product mix, and equipment availability [3]. Common areas that become bottlenecks in a wafer fab are the photolithography [20], ion implantation [21], and diffusion [18, 19] areas. This paper focuses on the diffusion area where the bottlenecks appear due the batching nature of the process to overcome the long processing times.
@@
For $\beta$, the focus is on articles that deal with multiple incompatible families (i.e., $P|p-batch,incompatible|\gamma$), although key articles with a single compatible family (i.e., $P|p-batch|\gamma$) are also considered, as the paper extends some of their concepts to the incompatible case. The review also includes some articles that consider non-identical job sizes ($s_{j}$) and release times ($r_{j}$). For $\gamma$, the reviewed objective functions primarily fall into three categories: (i) utilization objectives such as minimizing the makespan ($C_{\max}$); (ii) cycle time objectives such as minimizing the total completion time (TCT) and its weighted counterpart (TWCT); and (iii) on-time delivery objectives such as minimizing earliness, lateness, tardiness, number of tardy jobs, including their weighted variants [2]. At bottleneck areas, fab leaders often focus on two key aspects: the cycle time of the jobs to increase the effective movement of products [19] and the utilization of the machines. This paper focuses on the diffusion area using two objectives: the TWCT, which specifically targets the cycle time performance [2], and the makespan ($C_{\max}$), which specifically targets the utilization performance [2].
@@ VI-A The Automaton model
The Automaton (AU) model derives its name from the automaton global constraint it relies on, which is available in different open-source CP solvers [43, 24, 44]. Automaton constraints model finite state machines, enabling the representation of feasible transitions between states. In the AU model, these finite states represent the elapsed time when processing a batch of each family. Consider the automaton depicted in Figure 2, where nodes are states and arcs are transition possibilities. The descriptions in each state help understand the figure and the numbers in parenthesis ultimately identify them. Each arc is identified by a triplet $(i,k,\tau)$, where $i$ is the initial state, $k$ is the end state, and $\tau$ is the required transition value to trigger such transition from $i$ to $k$. For example, if a machine is empty and receives a transition value of 0, then the machine remains empty. But if it receives a transition value of 1, it means that a new batch of family 1 starts on the machine. From this point, the only possible transition values that the automaton can receive are the ones necessary to progress the states until completing the processing time $\delta^{1}$ of the batch. Once the batch is completed, three possible transitions can be made: either the machine becomes empty (0), or a new batch of family 1 starts (1), or a new batch of family 2 starts ($\delta^{1}+1$). From that point, a new sequence of transition values must be given such that the state transitions respect the automaton structure.
@@ VI-B The Assign-and-Schedule Model
Objective function (3) minimizes the TWCT. Constraints (3) assign each job to exactly one possible batch. Constraints (3) deactivate the presence of the batches if no jobs are assigned to them. Constraints (3) activate the presence of the batches if at least one job is assigned to them. Constraints (3) ensure that if a batch is used, its size respects the maximum batch capacity. Constraints (3) guarantee that batches are scheduled once all the jobs in them are available. Constraints (3) schedule each batch on exactly one machine, if present. Finally, constraints (3) use the global constraint $\texttt{noOverlap}{(\varphi)}$, which receives a sequence variable $\varphi$. This constraint ensures that the present interval variables in the provided sequence not overlap. Hence, constraints (3) guarantee that only one batch is processed at a time on each machine.
@@ VI-C The Synchronized Model
Objective function (4) minimizes the TWCT. Constraints (4) assign each job to exactly one batch on exactly one machine. Constraints (4) allow each batch to be processed in at most one machine. Constraints (4) use the $\texttt{synchronize}{(v,B)}$ global constraint, which receives an interval variable $v$ and a set of interval variables $B$. If $v$ is present, this constraint attempts to synchronize it with any intervals in $B$ that are also present [42], i.e., $\texttt{presenceOf}(v)=1\implies\texttt{startOf}(v)=\texttt{startOf}(b)\wedge\texttt{endOf}(v)=\texttt{endOf}(b),\forall\leavevmode\nobreak\ b\in B$ s.t. $\texttt{presenceOf}(b)=1$. In practice, when the interval variable representing the job being assigned to a batch on a machine (i.e., $x_{jm}^{b}$) is part of the solution, the synchronize constraint attempts to align its start and end times with those of the batch on the machine (i.e., $y_{bm}$), if it is also present. However, the synchronize constraint only ensures synchronization if both interval variables are present; it does not inherently enforce the presence of $y_{bm}$ when $x_{jm}^{b}$ is scheduled. This subtle issue, where the presence of $x_{jm}^{b}$ should logically imply the presence of $y_{bm}$ is addressed in the original model by Cakici et al. [10] by means of an additional constraints that defines the correct duration of the batches depending on the jobs assigned to it, if any. This constraint is necessary in their AM application as the duration of the batch depends on the size of the parts 3D-printed in it. However, this constraint is not required in the context of the diffusion area studied in this paper since the duration of the batches is already predefined. Nonetheless, the presence of the correct batches is still needed to be enforced. To resolve this, constraints (4) introduce an explicit implication constraint which ensures that, whenever $x_{jm}^{b}$ is included in the solution, the corresponding batch interval $y_{bm}$ is also present. Constraints (4) use the noOverlap global constraint to ensure that the present interval variables in the machine sequence do not overlap, guaranteeing at most one batch being processed at a time on each machine. Finally, constraints (4) guarantee that the load of the batches respect the maximum capacity.
@@ VII Computational experiments and results
and then setting $r_{j}\sim U\set{[C_{\max,lb}]}$. Finally, the maximum batch size for each family follows the distribution $u_{f}\sim U\{\Omega_{f}\}$, where $\Omega_{f}=\{u\in\Omega:u\geq\max_{j\in\mathcal{J}_{f}}n_{j}\}$ and $\Omega=\{10,15,20,25,30,35,40,45,50,75,100,125,150\}$. This approach ensures that instances are feasible by construction, allowing maximum batch sizes for each family that are larger than the size of any of its jobs.
@@
Table II summarizes the models implemented in each solver. The A model was only implemented in CP Optimizer because its state functions, alwaysEqual global constraint, and alignment parameters are only available in it. On the other hand, the AU model was only implemented in CP-SAT because the automaton global constraint required is not available in CP Optimizer (although it can be implemented in Choco-solver [24] and MiniZinc [44]). The AS, S, and RS models (and their variants with SB constraints) are implemented in both solvers.
@@
Figures 69 summarize our results. Figures 6 and 7 focus on the results for the TWCT, while Figures 8 and 9 focus on the $C_{\max}$ objective. Each graph presents the following characteristics:
@@
    Average relative gap (%) with its 95% confidence interval (CI).
@@
The relative gap for each instance is computed as follows. First, the minimum objective value $z_{i,obj}^{*}$ for each instance $i$ and objective $obj\in\{TWCT,C_{\max}\}$ is computed from the ones obtained by all the models in all the solvers. Then, for a given model and solver, the relative gap of the instance is defined as $\frac{|z_{i,obj,\text{model, solver}}-z_{i,obj}^{*}|}{|z_{i,obj}^{*}|}$. It is relative because it is with respect to the best known objective function value instead of the optimal value.
@@ VII-A Results with the TWCT objective
Figure 6(b) presents the average relative gap (%) along with its 95% confidence interval for each model in CP Optimizer. Although the A model did not prove optimality within the time limit in any instance, it consistently finds the best solutions, evident by the constant 0% relative gap of the black series. The RS and RS+SB models also obtained relative gaps close to 0%, and this is why their series are presented under the black one. In contrast, the AS and AS+SB models exhibit the worst average gaps, with wider confidence intervals. This is partly due to their failure to find solutions in several instances, resulting in fewer data points for averaging. Notably, a series break for these models appears at 200 jobs, 6 families, and 10 machines, reflecting a complete failure to solve any instance in that class. Furthermore, the consistently lower gaps of the AS+SB model compared to the AS model underscore the beneficial impact of incorporating SB constraints in the search process.
@@
Figure 7 focuses on the results of CP-SAT, which combines CP methods and satisfiability (SAT) methods in the search process. SAT methods benefit from binary structures in the problem. For this reason, in this solver, the AS model is able to find solutions on all the instances, as evidenced in Figure 7(a). The distinction between assignment and scheduling decisions of the AS model provides the proper structure for both CP and SAT methods to properly work, allowing the AS model to find solutions in all the instances; and even declare optimality in one of them. In contrast, the RS+SB model is unable to find solutions in most of the largest instances, precisely because of the additional SB constraints that become harder to satisfy within the time limit. The RS model is unable to find solutions in the largest instances with 200 jobs, precisely because of the amount of job-in-batch decisions that need to be made. On the other hand, the AU model clearly shows seasonal behavior around 10 machines.
When examining the average relative gaps in CP-SAT from Figure 7(b), it becomes evident that no model consistently matches the best results obtained by the A model in CP Optimizer. Nonetheless, the gaps are very close. The AU gap increases with the size of the instances. This is because of the $\tau_{mt}$ variables that represent the transition values provided to the automaton at every time step of the scheduling horizon, which increases with the number of jobs. Additionally, the element constraints set the appropriate transition values at the start times of the jobs on each machine. However, they do so only if the jobs are actually processed on such machine. While this could be seen as a reified constraint, it is not. Instead, this is indirectly achieved with the modified start times $\hat{s}_{jm}$ and modified transition values $\hat{\tau}_{jm}$. When the instances grow, more of these modified decisions have to be made, negatively impacting the performance of these element constraints, and causing the increase on the average relative gap.
@@
The best relative gaps in CP-SAT belong to the AS model, followed by the RS model, very close to 0%. Interestingly, the very same reason that makes The AS model model fall back in CP Optimizer, makes it shine in CP-SAT instead. The clear distinction between assignment and scheduling logic is well handled in CP-SAT thanks to the SAT component of this solver. Additionally, the RS model consistently produces comparable gaps of the AS model. However, as seen in Figure 7(a), the RS model is not able to find solutions in all the instances.
@@ VIII Concluding remarks
  • The A model consistently produces high-quality solutions (0% relative gap) but is limited by its reliance on proprietary solver features only available in the commercial solver CP Optimizer.
  • The clear separation between binary assignment and batch scheduling decisions enables the AS model to excel in the open-source solver CP-SAT. Thanks to the satisfiability components of the solver, which efficiently handle binary structures, the AS model achieves very small gaps and even proves optimality in isolated cases.